OpenStudy (anonymous):

It takes a lion 4 hours to eat a sheep. A leopard needs 5 hours and a wild dog needs 6 hours to eat a sheep. How long will it take the three of them together to eat a Sheep?

OpenStudy (anonymous):

3/4 hrs only

OpenStudy (anonymous):

show the workings

OpenStudy (anonymous):

Let x be the elapsed time for all three animals to eat one sheep. Solve the following equation for x:\[\frac{1}{4}+\frac{1}{5}+\frac{1}{6}=\frac{1}{x}\]\[x=\frac{6}{5}=1.2=1:12 \text{ hours}\]